# HAS-RRT: RRT-based Motion Planning using Topological Guidance

### v2

@@ I Introduction
This work presents the Hierarchical Annotated Skeleton RRT (HAS-RRT) method to quickly find workspace-guided paths in a single-query environment. Similar to [6], HAS-RRT  uses the workspace skeleton to search for paths. However, HAS-RRT  initially relies on the connectivity of the workspace to make large steps in its exploration and only retracts to focused local exploration when necessary. This approach makes HAS-RRT more robust to poor-quality skeletons than previous methods. Additionally, because the workspace skeleton can be annotated with information such as environment clearance or priority through certain passages, the strategy can be adapted to fit different motion planning applications. When the workspace skeleton contains workspace paths that can be extended to $\mathcal{C}$ paths, HAS-RRT exploits the guidance fully, and when the skeleton is not perfectly indicative of the motion planning solution, the method can use directional hints from such a skeleton and find a solution with additional exploration more gracefully than the other methods. This work is most applicable for robots whose workspace and configuration space are closely related, such as mobile robots.
@@ II-B Workspace-Guided Planning
Workspace skeletons are more effective when $\mathcal{C}$ is a subset of the workspace. When sampling along a workspace skeleton, a sample directly corresponds to a point in $\mathcal{C}$. When $\mathcal{C}$ does not correspond to the workspace (e.g. manipulator robots, where $\mathcal{C}$ has much different topology from the workspace), it can difficult to find a transformation from a sample taken in the workspace to a point in the free configuration space. EET [7] addresses this by corresponding a point in the the workspace to the location of the end effector, however, this approach does not guarantee finding a path if one exists.
@@ III-A The Algorithm
Expanding the tree (Lines 5 to 16 in Algorithm 1): These steps constitute the main loop that terminates when the query is solved or the algorithm runs out of resources. First, an active region is selected (see Section III-B for region selection details), and the planner samples inside for a direction to expand the tree toward. Once a valid $q_{rand}$ is sampled, an extension attempt is performed to expand the tree (Line 9 of Algorithm 1). The successful extension adds a new vertex to the tree and pushes the region that gave $q_{rand}$ to the end of its current skeleton edge. The region’s weight is also updated by incrementing its success record to increase its chances of being selected in the next iterations. A successful direct extension to the end of a skeleton edge and its branching into two regions for each adjacent edge are shown in Figures 3(c) and 3(d).
@@
Region Selection. The planner chooses the next region using a probability distribution based on prior extension success, the directed skeleton, and the explore/exploit bias. A candidate region $r$ is selected with probability $p_{r}$ calculated as follows, where $e$ is the explore bias hyperparameter and $R$ is the set of all regions:
@@ IV Validation
(a) 2D narrow passage and Medial Axis skeleton. We use a small 3-DOF rectangular prism robot. The robot at its widest is 60% of the width of the narrow passage.
(b) Grid tunnels and skeleton constructed with method from Section IV-B. We use a small 6-DOF L-shaped robot. The L shape is approximately 8% of the width of one of the blocks.
(c) Extended Z-shaped tunnels and a curated mean curvature skeleton. We use a cube-shaped robot that cannot rotate (3-DOF). The robot’s width is approximately 10% of the width of the tunnel.
(d) Grid Mining, with shafts (grey) and drifts (green), and skeleton constructed with method from Section IV-B. We use a small 6-DOF L-shaped robot. The L shape is approximately 8% of the width of one of the blocks.
(e) 8x8 3D grid maze and curated mean curvature skeleton. We use a stick shaped robot that can rotate around any axis 6-DOF. The length and depth of the stick are approximately 2% and 9% of the length of the environment respectively.
@@
We evaluate the method in five maze/tunnel environments shown and described in Figure 4. Some environments use the mean curvature skeleton, which may be challenging to HAS-RRT.
In Section IV-A and IV-B we describe our experimental setup and our novel modular way to construct a workspace skeleton for two of our environments. In Section IV-C we discuss the results of our experiments. Finally, in Section IV-D we analyze the performance of HAS-RRT with varying qualities of guidance to demonstrate how HAS-RRT behaves with poor guidance.
@@ IV-A Experimental Setup
The experiments were executed on a Google Cloud Compute e2-standard-4 computer with 4vCPUs and 16GB of RAM, running Ubuntu 20.04. All methods were implemented in our C++ Parasol Planning Library (PPL) [22] at the University of Illinois at Urbana-Champaign. Validation was done with the PQP-SOLID collision detection method [23]. Hyperparameters were chosen through experimentation.
@@ IV-B Skeleton Construction and Annotation
Instead of constructing an environment and computing a skeleton afterward, we simultaneously build an environment and skeleton. For simplicity, we create an environment with $90^{\circ}$ angles using cubic structures as building blocks. Considering orientation, there are $2^{6}$ ways to construct a cube with varying tunnels leading outwards. After designing the blocks, a skeleton is constructed for each block based on the tunnel openings, as shown in Figure 5(a). When blocks are paired, their respective skeletons are joined to make a connected graph representing the new topology (Figure 5(b)).
@@ IV-C Discussion
EET’s path costs are comparable and competitive to HAS-RRT because the Wavefront expansion guarantees any path to be centered in an environment. However, EET’s runtime is much higher than the other methods because the wavefront expansion is built probabilistically and thus must be recomputed for each seed. Additional investigation shows that the wavefront expansion’s build time takes up most of the runtime: build times for the Generated Grid environment averaged 66.49% of the total runtime and for the Extended Z Tunnel averaged 85.75%. The other two environments averaged similarly.
@@ IV-D Adapting to Skeleton Quality - Study
We used the same query used in the previous experiments from Section IV. Each vertex of the skeleton not corresponding to the query start and goal positions are shifted by a distance $d$ in a random direction and are guaranteed to be within the free workspace (see Figure 8). We increase the maximum radius of a sampling region from the previous set of experiments to accommodate the poorer-quality skeletons. Each experiment is run with 10 seeds.

### v3

@@ I Introduction
This work presents the Hierarchical Annotated Skeleton RRT (HAS-RRT) method to quickly find workspace-guided paths in a single-query environment. Similar to [6], HAS-RRT  uses the workspace skeleton to search for paths. However, HAS-RRT  initially relies on the connectivity of the workspace to make large steps in its exploration and only retracts to focused local exploration when necessary. This approach makes HAS-RRT more robust to poor-quality skeletons than previous methods. such as environment clearance or priority through certain passages , the strategy can be adapted to fit different motion planning applications. When the workspace skeleton contains workspace paths that can be extended to $\mathcal{C}$ paths, HAS-RRT exploits the guidance fully, and when the skeleton is not perfectly indicative of the motion planning solution, the method can use directional hints from such a skeleton and find a solution with additional exploration more gracefully than the other methods. This work most applies to robots whose workspace and configuration space are closely related, such as mobile robots.
@@ II-B Workspace-Guided Planning
Workspace skeletons are more effective when $\mathcal{C}$ is a subset of the workspace. When sampling along a workspace skeleton, a sample directly corresponds to a point in $\mathcal{C}$. When $\mathcal{C}$ does not correspond to the workspace (e.g. manipulator robots, where $\mathcal{C}$ has much different topology from the workspace), it can be difficult to find a transformation from a sample taken in the workspace to a point in the free configuration space. EET [7] addresses this by corresponding a point in the the workspace to the location of the end effector, however, this approach does not guarantee finding a path if one exists.
@@ III-A The Algorithm
Expanding the tree (Lines 5 to 16 in Algorithm 1): These steps constitute the main loop that terminates when the query is solved or the algorithm runs out of resources. First, an active region is selected (see Section III-B for region selection details), and the planner samples inside for a direction to expand the tree toward. Once a valid $q_{rand}$ is sampled, an extension attempt is performed to expand the tree (Line 9 of Algorithm 1). The successful extension adds a new vertex to the tree and pushes the region that gave $q_{rand}$ to the end of its current skeleton edge. The region’s weight is also updated by incrementing its success record to increase its chances of being selected in the next iterations. Then, new regions are created at the start of all outgoing new skeleton edges. A successful direct extension to the end of a skeleton edge and its branching into two regions for each adjacent edge are shown in Figures 3(c) and 3(d).
@@
Region Selection. The planner chooses the next region using a probability distribution based on prior extension success, the directed skeleton, and the explore/exploit bias.
A candidate region $r$ is selected with probability $p_{r}$ calculated as follows, where $e$ is the explore bias hyperparameter and $R$ is the set of all regions:
@@ IV Validation
(a) 2D narrow passage and Medial Axis skeleton. We use a small 3-DOF rectangular prism robot. The robot at its widest is 60% of narrow passage’s width.
(b) Grid tunnels and skeleton constructed with method from Section IV-B. We use a small 6-DOF L-shaped robot. The L shape is approximately 8% of a block’s width.
(c) Extended Z-shaped tunnels and a curated mean curvature skeleton. We use a cube-shaped robot that cannot rotate (3-DOF). The robot’s width is approximately 10% of the tunnel’s width.
(d) Grid Mining, with shafts (grey) and drifts (green), and skeleton constructed with method from Section IV-B. We use a small 6-DOF L-shaped robot. The L shape is approximately 8% of one block’s width.
(e) 8x8 3D grid maze and curated mean curvature skeleton. We use a stick shaped robot that can rotate around any axis 6-DOF. The length and depth of the stick are approximately 2% and 9% of the environment’s length, respectively.
@@
We evaluate the method in five maze/tunnel environments shown and described in Figure 4. In Section IV-A and IV-B we describe our experimental setup and our novel modular way to construct a workspace skeleton for two of our environments. In Section IV-C we discuss the results of our experiments. Finally, in Section IV-D we analyze the performance of HAS-RRT with varying qualities of guidance to demonstrate how HAS-RRT behaves with poor guidance.
@@ IV-A Experimental Setup
The experiments were executed on a Google Cloud Compute e2-standard-4 computer with 4vCPUs and 16GB of RAM, running Ubuntu 20.04. All methods were implemented in our C++ Parasol Planning Library (PPL) [22]. Validation was done with PQP-SOLID collision detection [23]. Hyperparameters were chosen through experimentation.
@@ IV-B Skeleton Construction and Annotation
In some cases, instead of constructing an environment and computing a skeleton afterward, we simultaneously build an environment and skeleton. For simplicity, we create an environment with $90^{\circ}$ angles using cubic structures as building blocks. Considering orientation, there are $2^{6}$ ways to construct a cube with varying tunnels leading outwards. After designing the blocks, a skeleton is constructed for each block based on the tunnel openings, as shown in Figure 5(a). When blocks are paired, their respective skeletons are joined to make a connected graph representing the new topology (Figure 5(b)).
@@ IV-C Discussion
EET’s path costs are comparable and competitive to HAS-RRT because the Wavefront expansion guarantees any path to be centered in an environment. However, EET’s runtime is much higher than the other methods because the wavefront expansion is built probabilistically and thus must be recomputed for each seed. Additional investigation shows that building the wavefront expansion takes up most of the runtime: build times for the Generated Grid environment averaged 66.49% of the total runtime and for the Extended Z Tunnel averaged 85.75%. The other two environments averaged similarly.
@@ IV-D Adapting to Skeleton Quality - Study
We used the same query used in the experiments from Section IV. Each vertex of the skeleton not corresponding to the query start and goal positions are shifted by a distance $d$ in a random direction and are guaranteed to be within the free workspace (see Figure 8). We increase the maximum radius of a sampling region from the previous set of experiments to accommodate the poorer-quality skeletons. Each experiment is run with 10 seeds.